Paul Coddington

Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244, U.S.A

Phone: (315) 443-4883 FAX: (315) 443-1973 Email: paulc@npac.syr.edu

October 20, 1994

Selection Committee Ref:C94/245 University of Ulster

To whom it may concern,

I was in the same group as James Allwright at Southampton University, and knew something of his work although we did not collaborate on any projects. Dr Allwright has worked with me on a couple of projects at the Northeast Parallel Architectures Center at Syracuse University.

Dr Allwright is very bright and strong academically, although he is somewhat lacking in communication skills. I collaborated with him on a research project on parallel graph coloring algorithms and he did very good work. He has a good working knowledge of parallel computing and parallel algorithms, and strong programming skills.

I highly recommend Dr Allwright for this position, as I believe he is well qualified, and would perform very well as a member of your software engineering team.

Yours sincerely,

Paul Coddington